OpenStudy (anonymous):

What is the solution to the equation log 4 x – log 4 2 = 2 ? x = 16 x = 4 x = 32 x = 8

OpenStudy (cwrw238):

log 4 x – log 4 2 = 2 by law of logs left hand side = log 4 x/2 log4 x/2 = 2 can you continue from here?

OpenStudy (cwrw238):

by definition if log to base a of b = c then a^c = b

OpenStudy (anonymous):

I am = still confused

OpenStudy (cwrw238):

ok we have log to base 4 of x/2 = 2 compare this with log to base a of b = c - so in your case you have a = 4 , b = x/2 and what is c = to?

OpenStudy (cwrw238):

c = 2 - do you see why? - i am comparing the two above formula formulae

OpenStudy (cwrw238):

* two above formulae

OpenStudy (cwrw238):

do you follow that?